zarkam21:

Could I multiply this by 6?

HuskyNation:

Full Question Please

zarkam21:

1 attachment
Vocaloid:

hint: how can you get the coefficients of y to be equal?

HuskyNation:

What is the point of multiply by 6? I don't see any question saying multiply

HuskyNation:

nvm i see

zarkam21:

1

Vocaloid:

hint: multiply the first equation by 2 and let me know what you get

zarkam21:

6x-2y=0

Vocaloid:

awesome, so you would write the new system of equations, replacing equation 1 by this new equation

Vocaloid:

new system: 6x - 2y = 0 5x + 2y = 22

Vocaloid:

and that's it for part I

zarkam21:

okay now for part II

Vocaloid:

for part II add the two equations together and you will be able to eliminate y.

zarkam21:

11x=22

Vocaloid:

good that's it for p II

Vocaloid:

for part III just solve for x

zarkam21:

x=2

Vocaloid:

awesome that's it for III for IV just substitute x = 2 into one of the original equations and solve for y

Vocaloid:

3x - y = 0 let x = 2 and solve for y.

zarkam21:

y=6

zarkam21:

oh I used the second equation to solve for y :S

Vocaloid:

good so your solution is (2,6)

Vocaloid:

it doesn't matter you can use either equation